OpenStudy (anonymous):

Describe how to transform the quantity of the fifth root of x to the seventh power, to the third powerinto an expression with a rational exponent Please someone help with this question ill give a medal

OpenStudy (anonymous):

\[\sqrt[5]{x}=x^{1/5}\] Raising to the 35th power, you have \[\left(x^{1/5}\right)^{35}=\left(\left(x^{1/5}\right)^5\right)^7=x^7\] Not sure if that's what you're looking for or not.